### 31 Mar 2026 v1

@@ 5.1 Base Encoder: ViT-MAE
STRADAViT builds on the ViT-MAE formulation (He et al., 2022), which uses a Vision Transformer (ViT) encoder (Dosovitskiy et al., 2021; Vaswani et al., 2017) with a lightweight decoder during masked reconstruction. Images are split into non-overlapping $p\times p$ patches and linearly embedded into a sequence of token vectors. A learnable class token is prepended and positional encodings are added, after which the sequence is processed by a stack of Transformer encoder blocks (multi-head self-attention and a feed-forward MLP).
Unless stated otherwise, we use the standard ViT-MAE base configuration (ViT-B/16). We do not modify the encoder embedding dimension, depth, attention head count, or the MAE decoder architecture relative to this default; controlled architectural comparisons are restricted to the optional projection head (contrastive objective only) and optional register tokens (below). We choose this MAE-family starting point because it provides a single, readily modifiable encoder family for both reconstruction and contrastive branches, allowing a controlled study of the transfer-learning training recipe without introducing additional implementation and compute variance from a second pretrained stack. All training views are resized to the model input resolution ($224\times 224$).
@@ 5.3 Register Tokens
We also evaluate an optional register-token variant inspired by recent foundation-style ViT designs (Oquab et al., 2023). In this setting, we insert $R$ learnable register tokens immediately after the class token and before the patch tokens. These register tokens participate in all encoder blocks and can act as additional content-bearing slots, while leaving the patch token grid unchanged. For compatibility with masked reconstruction and for consistent downstream readout, register tokens are removed from the sequence before the reconstruction decoder and before patch-token pooling in the contrastive objective. We report results for $R\in\{0,4\}$, where $R=0$ recovers the standard ViT-MAE encoder.
@@ 6.1 View Generation for Masked Reconstruction
After the geometric step, we apply the mild augmentation regime in Table 3, clamp to $[0,1]$, and then apply the initialization-dependent channel normalization described in Section 8.1. Figure 1 illustrates representative reconstruction-branch views produced by this pipeline.
@@ 6.2 Reconstruction Loss Variants
The reconstruction branch follows the ViT-MAE masked image modeling objective (He et al., 2022). Let $x$ denote an input view, $m$ a binary mask over patch locations, and $\hat{x}$ the model reconstruction (after unpatchifying decoder outputs). The default ViT-MAE training loss is a masked-patch mean squared error (MSE), computed only over the masked patches:
@@
We treat this as our baseline and compare two explicit modifications that add an additional reconstruction regularizer computed in the same normalized input space as $x$. In the following, $\langle\cdot\rangle$ denotes the mean over all pixels (and channels) of an image.
6.2.1 Baseline: Masked-Patch MSE (ViT-MAE Default)
The baseline objective is
@@ 6.2.1 Baseline: Masked-Patch MSE (ViT-MAE Default)
with weight $w_{\mathrm{L2}}=1.0$.
@@ 7 Contrastive Learning Objective
The contrastive objective is used in two settings: directly in the contrastive-only branch, and as the second stage of the two-stage branch. It learns invariances across multiple augmented views of the same standardized cutout. We apply the contrastive objective to the $\ell_{2}$-normalized projection-head outputs (Section 5), using all other views in the (distributed) batch as negatives. We compare three contrastive loss variants: a standard InfoNCE formulation, a “soft” hard-negative reweighting variant (soft-HCL), and a hard-negative objective (HCL) (Robinson et al., 2021).
@@ 7.1 View Generation for Contrastive Learning
We use IoA rather than a symmetric IoU criterion so that each additional view lies mostly inside the wide view while the wide view can retain broader context. Augmentations follow the same family as in the reconstruction branch, with the strong regime applied to the additional view. Each view is clamped to $[0,1]$ and normalized using the policy in Section 8.1.
@@ 7.2 Contrastive Loss Variants
Each training step operates on a batch of $B$ cutouts, generating $V$ views per cutout (in this work $V=2$) and producing an embedding $z\in\mathbb{R}^{d}$ for each view via the encoder, mean pooling over patch tokens, and the projection head (Section 5). Embeddings are $\ell_{2}$-normalized and concatenated across all devices to form a global set of keys; for each query embedding, positives are the other view(s) of the same cutout and negatives are all remaining keys (except the query itself). Similarities are computed as dot products and scaled by a temperature $\tau$ that is linearly warmed from $0.20$ to $0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$. In our distributed implementation, embeddings are all-gathered across devices to form the key set, but embeddings from other devices are treated as stop-gradient keys (i.e., gathered without gradient) while the loss is evaluated for the local queries.
The size of the negative pool is bounded by the distributed micro-batch used to compute the loss (i.e., a single forward pass). Let $G$ denote the number of devices (world size). Each loss evaluation involves $G\!\cdot\!B\!\cdot\!V$ key embeddings (all views across devices). For a given query, the number of negative keys is therefore approximately $G\!\cdot\!B\!\cdot\!V-V$ (excluding the $V$ views of the query’s underlying cutout, including itself). In the runs reported in this work we use $G=4$ GPUs, $B=64$ images per GPU, and $V=2$ views, yielding 256 images (512 keys) per micro-batch and 510 negative keys per query (equivalently, 255 other images). We use gradient accumulation to obtain a fixed effective global batch size of 2048 images per optimizer update. However, gradient accumulation does not increase the negative pool per loss evaluation. InfoNCE, HCL and soft-HCL share the same negative pool; they differ only in how negatives are weighted.
@@ 7.2.1 Baseline: InfoNCE
We use a standard temperature-scaled InfoNCE contrastive loss (the NT-Xent form used in SimCLR (Chen et al., 2020)) as our baseline. Let $s_{ij}=z_{i}^{\top}z_{j}$ denote the similarity between a query $z_{i}$ and a key $z_{j}$. Let $P(i)$ be the set of positive keys for query $i$ and $A(i)$ the set of allowed keys in the denominator (all keys excluding the query itself). Let $N(i)=A(i)\setminus P(i)$ denote the set of negatives. Define the positive mass
@@ 7.2.2 Hard Contrastive Learning (HCL)
We also evaluate hard-negative reweighting following Robinson et al. (2021). HCL keeps the same overall contrastive log-ratio structure as InfoNCE, but it modifies how the negative term is accumulated: negatives that are more similar to the query contribute more to the denominator.
@@ 8 Training Setup
This section summarizes (i) the initialization and input-normalization choices and (ii) the optimization settings and training schedules used in the reconstruction and contrastive branches.
8.1 Initialization and Input Normalization
We consider two starting points for continued pretraining: (i) continued pretraining from publicly available ViT-MAE weights pretrained on ImageNet-1K, and (ii) scratch training with random initialization. The choice of starting point also determines whether we apply an additional channel-wise input normalization step. When continuing from ImageNet-pretrained weights, we apply the standard ImageNet mean and standard deviation normalization to match the pretrained backbone. When training from scratch, we omit this normalization (identity), since our preprocessing already maps inputs to $[0,1]$. The contrastive-only branch initializes directly from the chosen starting checkpoint, whereas the two-stage branch initializes its contrastive stage from the corresponding reconstruction checkpoint trained with the same architectural settings.
8.2 Optimization and Training Hyperparameters
Unless stated otherwise, settings are shared across runs and the effective global batch size is fixed at 2048 images per optimizer update via gradient accumulation.
8.2.1 Common Optimization Settings
@@
8.2.2 Distributed Training and Normalization Details
@@
8.2.3 Reconstruction-Branch Training
Unless stated otherwise, the reconstruction branch trains for 35 epochs with a fixed mask ratio $0.75$ and patch size $16$. We use per-device batch size 256 and apply gradient clipping with $\lVert g\rVert_{2}$ capped at 1.0. We compare the three reconstruction losses described in Section 6 (masked-patch MSE; MSE + L1; MSE + brightness-weighted L1), using the weights specified there.
8.2.4 Contrastive-Branch Training
Unless stated otherwise, the contrastive branch trains for 35 epochs, both when used directly in the contrastive-only branch and as the second stage of the two-stage branch. We generate $V=2$ views per cutout and use per-device batch size 64 (so each step processes $2B$ views). The encoder is unmasked in this branch (no patch masking), and we apply gradient clipping with $\lVert g\rVert_{2}$ capped at 3.0. The contrastive temperature is warmed linearly from $\tau=0.20$ to $\tau=0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$.
@@ 9.1 Common Evaluation Setup
    Input normalization: Channel-wise normalization uses the mean and standard deviation stored with each checkpoint’s image processor: ImageNet mean/std for ImageNet-initialized checkpoints and identity normalization for scratch-trained checkpoints (Section 8.1).
@@ 10 Results and Analysis
Throughout, we report both linear probing and full fine-tuning because they probe different aspects of transfer. Linear probing remains a standard SSL diagnostic for frozen representation quality, and it is especially relevant when encoders are reused as generic backbones or when labeled data are limited. In our setting, improvements under a frozen-backbone probe can be larger and more consistent than improvements after end-to-end fine-tuning, where supervised adaptation can partially close gaps between pretraining variants. We therefore interpret probe gains as evidence of improved linear separability of the learned representations—a desirable property for transfer-oriented reuse—while treating fine-tuning as the primary measure of practical downstream performance on these benchmarks. All baselines and continued-pretraining variants are evaluated with the same downstream preprocessing, folds, and optimization protocol so that these comparisons remain controlled.
@@ 10.1 Baselines (DINOv2, ViT-MAE)
Under linear probing (Table 6), DINOv2 remains stronger than ViT-MAE. DINOv2-Base is the best baseline on all three datasets: Macro-F1 reaches $0.717\pm 0.014$ on MiraBest, $0.569\pm 0.005$ on LoTSS DR2, and $0.661\pm 0.004$ on RGZ DR1. DINOv2-Small remains consistently close behind. By contrast, the register-token DINOv2 variant is weaker on every benchmark in the frozen-feature setting, with the largest drop on RGZ DR1 ($0.661\pm 0.004\rightarrow 0.629\pm 0.001$). The strongest off-the-shelf starting point in terms of raw linear separability is therefore the non-register DINOv2 backbone.
Fine-tuning yields a less uniform ranking (Table 7). The strongest baseline is ViT-MAE-Base on MiraBest ($0.724\pm 0.002$), DINOv2-Base on LoTSS DR2 ($0.708\pm 0.008$), and DINOv2-Base(R) on RGZ DR1 ($0.812\pm 0.002$). On LoTSS DR2, DINOv2-Base is retained over DINOv2-Base(R) after a Macro-F1 tie at $0.708$ because it has the higher Weighted-F1. The register variant therefore no longer dominates the fine-tuning baselines, although it remains the strongest DINOv2 model on RGZ DR1. Optimization sensitivity also remains pronounced: the non-register DINOv2-Base variant is unstable on MiraBest ($0.461\pm 0.195$), whereas DINOv2-Small ($0.707\pm 0.031$), DINOv2-Base(R) ($0.704\pm 0.102$), and ViT-MAE-Base ($0.724\pm 0.002$) are more stable there. The fine-tuning ranking is therefore more sensitive to dataset and configuration than the linear-probe ranking.
@@ 10.2.1 Reconstruction-Only Branch
In this section we evaluate continued pretraining runs initialized from ImageNet-pretrained ViT-MAE weights (Section 8.1). These phase 1 results compare the reconstruction losses from Section 6 while keeping the rest of the protocol fixed.
@@ 10.2.2 Contrastive-Only Branch
We also evaluate the contrastive-only branch, in which the contrastive objective is applied directly from the chosen initialization checkpoint without a preceding reconstruction stage. We compare InfoNCE, Soft-HCL, and HCL under the same downstream protocol for the ViT-MAE-based runs. The separate DINOv2 initialization study is reported below under a fixed HCL recipe.
@@
The contrastive branch improves substantially on the phase 1 branch under both linear probing and fine-tuning. The main gain therefore comes from the contrastive objective rather than from masked reconstruction. Loss rankings remain dataset-dependent: Soft-HCL is strongest in several MiraBest and LoTSS DR2 settings, whereas HCL is the most reliable choice on RGZ DR1. Registers interact with the loss in the same dataset-dependent way and do not produce a uniform ordering across datasets.
@@ 10.2.3 DINOv2 Initialization Ablation
To test portability of the same adaptation pipeline to a differently pretrained ViT family, we also perform a targeted DINOv2 initialization ablation based on fixed HCL training. This provides a direct check of whether the radio astronomy adaptation pipeline transfers to a backbone whose original pretraining is already contrastive and more structured than ViT-MAE. Table 12 summarizes this ablation.
Table 12: DINOv2 initialization ablation under the contrastive-only HCL recipe (mean $\pm$ std over $K=3$ folds).
@@
Linear Probe DINOv2-Base
@@
Fine-Tuning DINOv2-Base
@@
The DINOv2-based results reported here show that the same contrastive adaptation recipe is not specific to the ViT-MAE starting point. Under linear probing, the DINOv2-Base(R)-initialized run improves on the corresponding off-the-shelf DINOv2-Base(R) baseline on all three datasets, with the largest change on RGZ DR1. Under full fine-tuning, the adapted model improves on the off-the-shelf DINOv2-Base(R) baseline on MiraBest and RGZ DR1, but not on LoTSS DR2. Relative to the ViT-MAE-based adapted models, the DINOv2-initialized run is broadly competitive, with only small differences across all three datasets. The main exception remains LoTSS DR2, where the adapted DINOv2 result follows the same pattern as the ViT-MAE-based adapted models and remains below the off-the-shelf baselines.

### 08 Apr 2026 v3

@@ 5.1 Base Encoder: ViT-MAE
STRADAViT builds on the ViT-MAE formulation (He2022MAE), which uses a Vision Transformer (ViT) encoder (Dosovitskiy2021ViT; Vaswani2017Attention) with a lightweight decoder during masked reconstruction. Images are split into non-overlapping $p\times p$ patches and linearly embedded into a sequence of token vectors, which is then processed by a stack of Transformer encoder blocks (multi-head self-attention and a feed-forward MLP).
Unless stated otherwise, we use the standard ViT-MAE base configuration (ViT-B/16). We do not modify the encoder embedding dimension, depth, attention head count, or the MAE decoder architecture relative to this default; controlled architectural comparisons are restricted to the optional projection head (contrastive objective only) and optional register tokens (below). We choose this MAE-family starting point because it provides a single, readily modifiable encoder family for both reconstruction and contrastive branches, allowing a controlled study of the transfer-learning training recipe without introducing additional implementation and compute variance from a second pretrained stack.
@@ 5.3 Register Tokens
We also evaluate an optional register-token variant inspired by recent foundation-style ViT designs (Oquab2023DINOv2). In this setting, we insert $R$ learnable register tokens immediately after the class token and before the patch tokens. These register tokens participate in all encoder blocks and can act as additional content-bearing slots, while leaving the patch token grid unchanged. For compatibility with masked reconstruction and for consistent downstream readout, register tokens are removed from the sequence before the reconstruction decoder and before patch-token pooling in the contrastive objective. We report results for $R\in\{0,4\}$, where $R=0$ recovers the standard ViT-MAE encoder.
@@ 6.1 View Generation for Masked Reconstruction
After the geometric step, we apply the mild augmentation regime in Table 3, clamp to $[0,1]$, and then apply model-dependent channel normalization (i.e., ImageNet normalization for pretrained backbones). Figure 1 illustrates representative reconstruction-branch views produced by this pipeline.
@@ 6.2 Reconstruction Loss Variants
The reconstruction branch follows the ViT-MAE masked image modeling objective (He2022MAE). Let $x$ denote an input view, $m$ a binary mask over patch locations, and $\hat{x}$ the model reconstruction (after unpatchifying decoder outputs). The default ViT-MAE training loss is a masked-patch mean squared error (MSE), computed only over the masked patches:
@@
We treat this as our baseline, denote it by $\mathcal{L}_{\mathrm{L2}}$, and set its weight to $w_{\mathrm{L2}}=1.0$:
@@
We then compare two explicit modifications that add an additional reconstruction regularizer computed in the same normalized input space as $x$. In the following, $\langle\cdot\rangle$ denotes the mean over all pixels (and channels) of an image.
@@ 7 Contrastive Learning Objective
The contrastive objective is used in two settings: directly in the contrastive-only branch, and as the second stage of the two-stage branch. It learns invariances across multiple augmented views of the same standardized cutout. We apply the contrastive objective to the $\ell_{2}$-normalized projection-head outputs (Section 5), using all other views in the (distributed) batch as negatives. We compare three contrastive loss variants: a standard InfoNCE formulation, a “soft” hard-negative reweighting variant (soft-HCL), and a hard-negative objective (HCL) (Robinson2021HardNegatives).
@@ 7.1 View Generation for Contrastive Learning
We use IoA rather than a symmetric IoU criterion so that each additional view lies mostly inside the wide view while the wide view can retain broader context. Augmentations follow the same family as in the reconstruction branch, with the strong regime applied to the additional view. Each view is clamped to $[0,1]$ and normalized using the same model-dependent policy.
@@ 7.2 Contrastive Loss Variants
Each training step operates on a batch of $B$ cutouts, generating $V$ views per cutout (in this work $V=2$) and producing an embedding $z\in\mathbb{R}^{d}$ for each view via the encoder, mean pooling over patch tokens, and the projection head (Section 5). Similarities are computed as dot products and scaled by a temperature $\tau$ that is linearly warmed from $0.20$ to $0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$. In our distributed implementation, the normalized embeddings are all-gathered across devices to form the key set, but embeddings from other devices are treated as stop-gradient keys (i.e., gathered without gradient) while the loss is evaluated for the local queries. In the ViT-MAE-based runs reported here, 4 GPUs with per-device batch size 64 and 2 views per cutout yield 512 keys per micro-batch and 510 negatives per query; gradient accumulation increases the effective optimizer batch size but does not enlarge this per-step negative pool. InfoNCE, HCL, and soft-HCL therefore share the same negative pool and differ only in how negatives are weighted.
@@ 7.2.1 Baseline: InfoNCE
We use a standard temperature-scaled InfoNCE contrastive loss (the NT-Xent form used in SimCLR (Chen2020SimCLR)) as our baseline. Let $s_{ij}=z_{i}^{\top}z_{j}$ denote the similarity between a query $z_{i}$ and a key $z_{j}$. Let $P(i)$ be the set of positive keys for query $i$ and $A(i)$ the set of allowed keys in the denominator (all keys excluding the query itself). Let $N(i)=A(i)\setminus P(i)$ denote the set of negatives. Define the positive mass
@@ 7.2.2 Hard Contrastive Learning (HCL)
We also evaluate hard-negative reweighting following Robinson2021HardNegatives. HCL keeps the same overall contrastive log-ratio structure as InfoNCE, but it modifies how the negative term is accumulated: negatives that are more similar to the query contribute more to the denominator.
@@ 8 Training Setup
This section summarizes the optimization settings and training schedules used in the reconstruction and contrastive branches. All runs use continued pretraining from publicly available ViT-MAE weights pretrained on ImageNet-1K, with standard ImageNet mean and standard deviation normalization applied to match the pretrained backbone. The contrastive-only branch initializes directly from the chosen starting checkpoint, whereas the two-stage branch initializes its contrastive stage from the corresponding reconstruction checkpoint trained with the same architectural settings. Unless stated otherwise, settings are shared across runs and the effective global batch size is fixed at 2048 images per optimizer update via gradient accumulation.
8.1 Common Optimization Settings
@@
8.2 Distributed Training and Normalization Details
@@
8.3 Reconstruction-Branch Training
The reconstruction branch trains for 35 epochs with a fixed mask ratio $0.75$ and patch size $16$. We use per-device batch size 256 and apply gradient clipping with $\lVert g\rVert_{2}$ capped at 1.0. We compare the three reconstruction losses described in Section 6 (masked-patch MSE; MSE + L1; MSE + brightness-weighted L1), using the weights specified there.
8.4 Contrastive-Branch Training
The contrastive branch trains for 35 epochs, both when used directly in the contrastive-only branch and as the second stage of the two-stage branch. We generate $V=2$ views per cutout and use per-device batch size 64 (so each step processes $2B$ views). The encoder is unmasked in this branch (no patch masking), and we apply gradient clipping with $\lVert g\rVert_{2}$ capped at 3.0. The contrastive temperature is warmed linearly from $\tau=0.20$ to $\tau=0.10$ over the warmup portion of training (first 15% of epochs) and then kept at $\tau=0.10$.
@@ 9.1 Common Evaluation Setup
    Input normalization: Channel-wise normalization uses the mean and standard deviation stored with each checkpoint’s image processor, i.e. ImageNet mean/std for ImageNet-initialized checkpoints.
@@ 10 Results and Analysis
Throughout, we report both linear probing and full fine-tuning because they probe different aspects of transfer. Linear probing remains a standard SSL diagnostic for frozen representation quality, and it is especially relevant when encoders are reused as generic backbones or when labeled data are limited. In our setting, improvements under a frozen-backbone probe can be larger and more consistent than improvements after end-to-end fine-tuning, where supervised adaptation can partially close gaps between pretraining variants. We therefore interpret probe gains as evidence of improved linear separability of the learned representations—a desirable property for transfer-oriented reuse—while treating fine-tuning as the primary measure of practical downstream performance on these benchmarks. In other words, linear probing serves as the sharper representation-quality diagnostic, whereas fine-tuning serves as the practical downstream check.
@@ 10.1 Baselines (DINOv2, ViT-MAE)
Under linear probing (Table 6), DINOv2 remains stronger than ViT-MAE. DINOv2-Base is the best baseline on all three datasets: Macro-F1 reaches $0.717\pm 0.014$ on MiraBest, $0.569\pm 0.005$ on LoTSS DR2, and $0.661\pm 0.004$ on RGZ DR1. DINOv2-Small remains consistently close behind. By contrast, the register-token DINOv2 variant is weaker on every benchmark in the frozen-feature setting, with the largest drop on RGZ DR1 ($0.661\pm 0.004\rightarrow 0.629\pm 0.001$). The strongest off-the-shelf baseline under linear probing is therefore the non-register DINOv2 backbone.
Fine-tuning yields a less uniform ranking (Table 7). The strongest baseline is ViT-MAE-Base on MiraBest ($0.724\pm 0.002$), DINOv2-Base on LoTSS DR2 ($0.708\pm 0.008$), and DINOv2-Base(R) on RGZ DR1 ($0.812\pm 0.002$). On LoTSS DR2, DINOv2-Base is retained over DINOv2-Base(R) after a Macro-F1 tie at $0.708$ because it has the higher Weighted-F1. Optimization sensitivity also remains pronounced: the non-register DINOv2-Base variant is unstable on MiraBest ($0.461\pm 0.195$), whereas DINOv2-Small ($0.707\pm 0.031$), DINOv2-Base(R) ($0.704\pm 0.102$), and ViT-MAE-Base ($0.724\pm 0.002$) are more stable there. The fine-tuning ranking is therefore more sensitive to dataset and configuration than the linear-probe ranking.
@@ 10.2.1 Reconstruction-Only Branch
In this section we evaluate continued pretraining runs initialized from ImageNet-pretrained ViT-MAE weights. These phase 1 results compare the reconstruction losses from Section 6 while keeping the rest of the protocol fixed.
@@ 10.2.2 Contrastive-Only Branch
We also evaluate the contrastive-only branch, in which the contrastive objective is applied directly from the chosen initialization checkpoint without a preceding reconstruction stage. We compare InfoNCE, Soft-HCL, and HCL under the same downstream protocol for the ViT-MAE-based runs.
@@
The contrastive branch yields substantial improvements over phase 1 under both linear probing and fine-tuning. The main gain therefore comes from the contrastive objective rather than from masked reconstruction. Loss rankings remain dataset-dependent: Soft-HCL is strongest in several MiraBest and LoTSS DR2 settings, whereas HCL is the most reliable choice on RGZ DR1. This pattern is consistent with differences in class structure and confusability across datasets, and registers interact with the loss in the same dataset-dependent way rather than producing a uniform ordering.
@@ 10.2.3 DINOv2 Initialization Ablation
To test portability of the same adaptation pipeline to a differently pretrained ViT family, we also perform a targeted DINOv2 initialization ablation based on fixed HCL training. This assesses whether the radio astronomy adaptation pipeline transfers to a backbone whose original pretraining is already contrastive and more structured than ViT-MAE. Table 12 summarizes this ablation.
Table 12: DINOv2 initialization ablation under the contrastive-only HCL recipe (mean $\pm$ std over $K=3$ folds). Best Macro-F1 per dataset within each transfer setting is highlighted in bold; ties are broken by Weighted-F1.
@@
Linear Probe DINOv2-Base $0.683\pm 0.010$ $0.683\pm 0.010$ $0.551\pm 0.010$ $0.724\pm 0.006$ $0.735\pm 0.004$ $0.894\pm 0.000$
@@
Fine-Tuning DINOv2-Base $0.721\pm 0.020$ $0.722\pm 0.020$ $0.676\pm 0.009$ $0.816\pm 0.004$ $0.816\pm 0.000$ $0.919\pm 0.000$
@@
The DINOv2-based results show that the same contrastive adaptation recipe is not specific to the ViT-MAE starting point. These results should be read as a constrained initialization study, because the DINOv2 runs use a smaller micro-batch and therefore a smaller negative pool than the ViT-MAE runs. Under the fixed HCL recipe, DINOv2-Base(R) is the stronger adapted initialization on all three datasets in both transfer settings, even though the off-the-shelf linear-probe baselines favored DINOv2-Base without registers. The adaptation effect therefore depends on initialization family and does not simply preserve the off-the-shelf ranking. Relative to their own off-the-shelf baselines, the adapted DINOv2 variants improve on MiraBest and RGZ DR1, whereas LoTSS DR2 remains below the off-the-shelf DINOv2 results. Relative to the ViT-MAE-based adapted models, the best DINOv2-initialized run is broadly competitive and marginally stronger than the ViT-MAE-based contrastive-only models on MiraBest and LoTSS DR2 under fine-tuning. The ablation therefore supports portability of the adaptation recipe.